# Selmer stability for elliptic curves in Galois $\ell$-extensions

### v1

@@ 2.3. The method of Reichardt and Scholz
$\chi^{(q)}:\operatorname{Gal}(\mathbb{Q}(\mu_{q})/\mathbb{Q})\rightarrow\mathbb{Z}/\ell\mathbb{Z}.$
Identifying $\operatorname{Gal}(\mathbb{Q}(\mu_{q})/\mathbb{Q})$ with $(\mathbb{Z}/q\mathbb{Z})^{\times}$, and the $\ell$-torsion in $(\mathbb{Z}/q\mathbb{Z})^{\times}$ with $\mathbb{Z}/\ell\mathbb{Z}$, $\chi^{(q)}$ is given by $\chi^{(q)}(a):=a^{(q-1)/\ell}$. Set $\mathbb{L}:=L\left(\mu_{\ell^{N}}\cup\{\sqrt[\ell]{p}|p\in X\}\right)$ and let $\Delta:=\operatorname{Gal}(\mathbb{L}/\mathbb{Q})$. There exists $\sigma\in\Delta$ such that, for any prime $q$ unramified in $\mathbb{L}$ with $\sigma_{q}=\sigma$, the following conditions hold:
@@
[…]
Such an element exists since $FL\cap\mathbb{Q}\left(\mu_{\ell^{N}}\cup\{\sqrt[\ell]{p}|p\in X\}\right)=\mathbb{Q}$ by Lemma 2.14. Let $\widetilde{\varphi}:\operatorname{G}_{\mathbb{Q}}\to\widetilde{G}$ correspond to $\widetilde{L}$, and define $\widetilde{\psi}:=\widetilde{\varphi}(\chi^{(q)})^{-1}$, with the corresponding extension denoted $\widetilde{L}^{\prime}/L/\mathbb{Q}$. This modified extension solves the embedding problem for $\widetilde{G}$ while ensuring that property $(\mathfrak{S}_{N})$ is satisfied.
@@
Set $N$ to be such that #$\#G=\ell^{N}$. Recall the filtration on $G$ by subgroups $G_{i}$ as in (2.1). By induction on $i$, we construct Galois extensions
@@ 3.1. Selmer groups and diophantine stability in cyclic $\ell$-extensions
We summarize and recall the results in [PR25, section 2]. Throughout this section, $\ell$ is a fixed odd rational prime number and $L/K$ a Galois extension of number fields with $\operatorname{Gal}(L/K)=G\simeq\mathbb{Z}/\ell\mathbb{Z}$. Let $M$ be a finite dimensional $\mathbb{F}_{\ell}$ vector space equipped with the structure of a $G$-module. Define $M^{G}:=\{m\in M\mid gm=m\text{ for all }g\in G\}$ and $M_{G}:=M/\langle(g-1)m\mid g\in G,m\in M\rangle$. The following lemma is useful.
@@
See [NSW08, Proposition 1.6.12]. ∎
Let $E$ be a fixed elliptic curve defined over $K$. Given a finite extension $F/K$, denote by $E(F)$ (resp. $\Sh(E/F)$) the Mordell–Weil group (resp. Tate–Shafarevich group) of $E$ over $F$. Fix an algebraic closure $\overline{L}$ of $L$ and let $\operatorname{G}_{L}=\operatorname{Gal}(\overline{L}/L)$ denote the absolute Galois group of $L$. Let $\Omega_{L}$ be the set of all non-archimedean places of $L$, and for each $v\in\Omega_{L}$, let $L_{v}$ denote the completion of $L$ at $v$. Choose an algebraic closure $\overline{L}_{v}$ of $L_{v}$ and fix an embedding $\iota_{v}:\overline{L}\hookrightarrow\overline{L}_{v}$. This choice induces an embedding of Galois groups $\iota_{v}^{*}:\operatorname{G}_{L_{v}}\hookrightarrow\operatorname{G}_{L}$, where we set $\operatorname{G}_{L_{v}}:=\operatorname{Gal}(\overline{L}_{v}/L_{v})$. For notational convenience, we write
$H^{i}(L,\cdot):=H^{i}(\operatorname{G}_{L},\cdot),\quad H^{i}(L_{v},\cdot):=H^{i}(\operatorname{G}_{L_{v}},\cdot)$
@@
$\kappa_{E,v}^{(n)}:\frac{E(L_{v})}{\ell^{n}E(L_{v})}\hookrightarrow H^{1}(L_{v},E[\ell^{n}])$
@@
$\kappa_{E,v}:E(L_{v})\otimes\left(\mathbb{Q}_{\ell}/\mathbb{Z}_{\ell}\right)\hookrightarrow H^{1}(L_{v},E[\ell^{\infty}]).$
@@
$\operatorname{Sel}_{\ell^{n}}(E/L):=\ker\left(H^{1}(L,E[\ell^{n}])\to\prod_{v\in\Omega_{L}}H^{1}(L_{v},E)[\ell^{n}]\right).$
@@
$0\to E(L)\otimes\mathbb{Q}_{\ell}/\mathbb{Z}_{\ell}\to\operatorname{Sel}_{\ell^{\infty}}(E/L)\to\Sh(E/L)[\ell^{\infty}]\to 0$
@@
$\operatorname{Sel}_{\ell}(E/L):=\ker\left(H^{1}(L_{S}/L,E[\ell])\xrightarrow{\Phi_{S,L}}\bigoplus_{w\in S(L)}H^{1}(L_{w},E)[\ell]\right),$
@@
cf. [CS10, p.8]. For a prime $v$ of $K$, define the restriction map
$\gamma_{v}:H^{1}(K_{v},E)[\ell]\to\bigoplus_{w\mid v}H^{1}(L_{w},E)[\ell],$
@@
$0\rightarrow E(K)\otimes\mathbb{Z}/\ell\mathbb{Z}\rightarrow\operatorname{Sel}_{\ell}(E/K)\rightarrow\Sh(E/K)[\ell]\rightarrow 0$
@@
It follows from the assumptions above that $\gamma_{\ell}$ is injective for every prime $\ell\in S$ and thus by Proposition 3.3, $\operatorname{Sel}_{\ell}(E/L)=0$. For further details, we refer to the proof of [PR25, Proposition 2.9]. ∎
Let $\rho_{E,\ell}:\operatorname{G}_{\mathbb{Q}}\to\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ be the Galois representation on $E[\ell]$, and let $\mathbb{Q}(E[\ell])$ be the field fixed by $\ker\rho_{E,\ell}$, so that $\operatorname{Gal}(\mathbb{Q}(E[\ell])/\mathbb{Q})\simeq\operatorname{im}\rho_{E,\ell}$. For a prime $p$ of good reduction for $E$, let $\operatorname{G}_{\mathbb{Q}_{p}}=\operatorname{Gal}(\overline{\mathbb{Q}}_{p}/\mathbb{Q}_{p})$ with inertia subgroup $\operatorname{I}_{p}$, and choose a lift $\sigma_{p}\in\operatorname{G}_{\mathbb{Q}_{p}}$ of Frobenius. The trace and determinant of $\rho_{E,\ell}(\sigma_{p})$ satisfy
@@
For $p\in\mathfrak{T}_{E,K}$, $\widetilde{E}(k_{v})[\ell]=0$ for any of the primes of $K$ such that $v|\ell$ (cf. [PR25, Lemma 2.11]).
@@
$\mathfrak{d}(\mathfrak{T}_{E,K})=\frac{\ell^{2}-\ell-1}{[K(\mu_{\ell}):\mathbb{Q}(\mu_{\ell})]\,(\ell^{2}-1)(\ell-1)}.$
@@ 3.2. Main results
$\varrho:\operatorname{Gal}\big{(}\mathbb{Q}(E[\ell])/\mathbb{Q}(\mu_{\ell})\big{)}\xrightarrow{\sim}\operatorname{SL}_{2}(\mathbb{F}_{\ell}).$
Set $F_{0}:=F\,\cap\,\mathbb{Q}(E[\ell])$ and
$N:=\varrho\left(\operatorname{Gal}(\mathbb{Q}(E[\ell])/F_{0})\right)\subseteq\operatorname{SL}_{2}(\mathbb{F}_{\ell}).$
@@
$[\operatorname{SL}_{2}(\mathbb{F}_{\ell}):N]=[F_{0}:\mathbb{Q}(\mu_{\ell})]=\ell^{k},$
for some $1\leq\ell^{k}\leq[F:\mathbb{Q}(\mu_{\ell})]$. For primes $\ell\geq 5$, Galois [Gal46, p. 412] showed that $\operatorname{PSL}_{2}(\mathbb{F}_{\ell})$ is simple. Therefore, $\operatorname{SL}_{2}(\mathbb{F}_{\ell})$ does not contain a proper normal subgroup $N$ with odd index $[\operatorname{SL}_{2}(\mathbb{F}_{\ell}):N]$. It thus follows that $\ell^{k}=1$ and $F_{0}=\mathbb{Q}(\mu_{\ell})$. ∎
@@
For ease of notation, we set $\mathfrak{T}_{E}:=\mathfrak{T}_{E,\mathbb{Q}}$. Let $\Sigma$ be a finite set of primes containing $\ell$ and the primes at which $E$ has bad reduction. The result below is used in conjunction with Proposition 3.4 is used to prove Theorem 3.10, which is the main result of this section.
@@
Let $G$ be an $\ell$-group with #$\#G=\ell^{n}$ and let $N\geq n$. There are infinitely many Galois extensions $L/\mathbb{Q}$ such that:
@@
For such a $p$, $p$ is totally ramified in $L^{p}$ and thus the inertial degree of $p$ in $L^{p}$ is $1$. Therefore, $L^{p}$ satisfies the conditions (i), (ii), (iii). We show that $p$ as above is determined by certain non-empty Chebotarev conditions. In other words, there exists a Galois extension $\mathcal{F}/\mathbb{Q}$ and a conjugation invariant non-empty subset $\mathcal{S}\subseteq\operatorname{Gal}(\mathcal{F}/\mathbb{Q})$, such that a prime $p$ which is unramified in $\mathcal{F}$ satisfies (1), (2), (3) if and only if the Frobenius element $\sigma_{p}\in\mathcal{S}$. The condition that $p\equiv 1\pmod{\ell^{N}}$ is equivalent to the requirement that $p$ splits completely in $\mathbb{Q}(\mu_{\ell^{N}})$, i.e., $\sigma_{p}=1$. The primes $q\in\Sigma$ are required to split in $L^{p}$. In other words, $q$ is an $\ell$-th power modulo $p$. Thus, $p$ is required to split completely in $\mathcal{F}^{\prime}:=\mathbb{Q}(\mu_{\ell^{N}}\cup\{\sqrt[\ell]{q}\mid q\in\Sigma\})$. On the other hand, by Lemma 3.8, $\mathcal{F}^{\prime}\cap\mathbb{Q}(E[\ell])=\mathbb{Q}(\mu_{\ell})$. Setting $\mathcal{F}:=\mathcal{F}^{\prime}\cdot\mathbb{Q}(E[\ell])=\mathbb{Q}(E[\ell]\cup\mu_{\ell^{N}}\cup\{\sqrt[\ell]{q}\mid q\in\Sigma\})$, one finds that
@@
Let $\mathcal{S}:=\mathcal{S}_{1}\times\mathcal{S}_{2}$ where $\mathcal{S}_{1}=\{1\}\subset\operatorname{Gal}(\mathcal{F}^{\prime}/\mathbb{Q}(\mu_{\ell}))$ and $\mathcal{S}_{2}\subset\operatorname{Gal}(\mathbb{Q}(E[\ell])/\mathbb{Q}(\mu_{\ell}))$ consisting of $\sigma$ such that $\operatorname{trace}\rho_{E,\ell}(\sigma)\not\equiv 2\pmod{\ell}$. It is clear that if $p\notin\Sigma$ is a prime which is unramified in $\mathcal{F}$ and $\sigma_{p}\in\mathcal{S}$, then the conditions (1)–(3) are satisfied. By the Chebotarev density theorem, this set of primes $p$ has positive density.
@@
These conditions are determined by non-empty Chebotarev sets. Condition (a) is determined in the field extension $\mathbb{Q}(E[\ell])/\mathbb{Q}(\mu_{\ell})$, and (b),(c) are determined in $\mathcal{F}^{\prime}:=L(\mu_{\ell^{N}},\sqrt[\ell]{v}\mid v\in\Sigma\cup\{p_{1},\dots,p_{m}\})$. Let $\mathcal{S}_{1}\subset\operatorname{Gal}(\mathcal{F}^{\prime}/\mathbb{Q}(\mu_{\ell}))$ be the Chebotarev set which corresponds to conditions (b) and (c), and $\mathcal{S}_{2}\subset\operatorname{Gal}(\mathbb{Q}(E[\ell])/\mathbb{Q}(\mu_{\ell}))$ correspond to condition (a). Note that $\mathcal{F}^{\prime}$ is an $\ell$-extension of $\mathbb{Q}(\mu_{\ell})$. Lemma 3.8 thus implies that $\mathcal{F}^{\prime}\cap\mathbb{Q}(E[\ell])=\mathbb{Q}(\mu_{\ell})$. Therefore,
@@
and let $\mathcal{S}:=\mathcal{S}_{1}\times\mathcal{S}_{2}$. By construction, $\mathcal{S}\subset\operatorname{Gal}(\mathcal{F}/\mathbb{Q})$ is the non-empty Chebotarev set which determines conditions (a)–(c). Thus, by the Chebotarev density theorem, the set of primes satisfying (a)–(c) has positive density. We choose one such prime $p$ and take $p_{m+1}:=p$. Let $\chi^{(p)}:\operatorname{Gal}(\mathbb{Q}(\mu_{p})/\mathbb{Q})\rightarrow\mathbb{Z}/\ell\mathbb{Z}$ be the associated character and take $\widetilde{L}$ be the twist of $\widetilde{L}_{0}$ by $(\chi^{(p)})^{-1}$ as in Definition 2.10. This modification satisfies all the required conditions and completes the inductive argument. ∎
@@
Let $\ell\geq 5$ be a prime and $E$ be an elliptic curve over $\mathbb{Q}$. Assume that $\rho_{E,\ell}:\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{GL}_{2}(\mathbb{F}_{\ell})$ is surjective and that $\operatorname{Sel}_{\ell}(E/\mathbb{Q})=0$. Let $G$ be an $\ell$-group. Then there are infinitely many $G$-extensions $L/\mathbb{Q}$ for which $\operatorname{Sel}_{\ell}(E/L)=0$.

### v2

@@ 2.3. The method of Reichardt and Scholz
$\chi^{(q)}\mathrel{\mathop{\ordinarycolon}}\operatorname{Gal}(\mathbb{Q}(\mu_{q})/\mathbb{Q})\rightarrow\mathbb{Z}/\ell\mathbb{Z}.$
Identifying $\operatorname{Gal}(\mathbb{Q}(\mu_{q})/\mathbb{Q})$ with $(\mathbb{Z}/q\mathbb{Z})^{\times}$, and the $\ell$-torsion in $(\mathbb{Z}/q\mathbb{Z})^{\times}$ with $\mathbb{Z}/\ell\mathbb{Z}$, $\chi^{(q)}$ is given by $\chi^{(q)}(a)\mathrel{\mathop{\ordinarycolon}}=a^{(q-1)/\ell}$. Set $\mathbb{L}\mathrel{\mathop{\ordinarycolon}}=L\left(\mu_{\ell^{N}}\cup\{\sqrt[\ell]{p}|p\in X\}\right)$ and let $\Delta\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\mathbb{L}/\mathbb{Q})$. There exists $\sigma\in\Delta$ such that, for any prime $q$ unramified in $\mathbb{L}$ with $\sigma_{q}=\sigma$, the following conditions hold:
@@
  • $\sigma(\sqrt[\ell]{p})=c_{p}\sqrt[\ell]{p}$ for all primes $p\in X$.
Such an element exists since $FL\cap\mathbb{Q}\left(\mu_{\ell^{N}}\cup\{\sqrt[\ell]{p}|p\in X\}\right)=\mathbb{Q}$ by Lemma 2.14. Let $\widetilde{\varphi}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\to\widetilde{G}$ correspond to $\widetilde{L}$, and define $\widetilde{\psi}\mathrel{\mathop{\ordinarycolon}}=\widetilde{\varphi}(\chi^{(q)})^{-1}$, with the corresponding extension denoted $\widetilde{L}^{\prime}/L/\mathbb{Q}$. This modified extension solves the embedding problem for $\widetilde{G}$ while ensuring that property $(\mathfrak{S}_{N})$ is satisfied.
@@
Set $N$ to be such that $G=\ell^{N}$. Recall the filtration on $G$ by subgroups $G_{i}$ as in (2.1). By induction on $i$, we construct Galois extensions
@@ 3.1. Selmer groups and diophantine stability in cyclic $\ell$-extensions
We summarize and recall the results in [pathakray, section 2]. Throughout this section, $\ell$ is a fixed odd rational prime number and $L/K$ a Galois extension of number fields with $\operatorname{Gal}(L/K)=G\simeq\mathbb{Z}/\ell\mathbb{Z}$. Let $M$ be a finite dimensional $\mathbb{F}_{\ell}$ vector space equipped with the structure of a $G$-module. Define $M^{G}\mathrel{\mathop{\ordinarycolon}}=\{m\in M\mid gm=m\text{ for all }g\in G\}$ and $M_{G}\mathrel{\mathop{\ordinarycolon}}=M/\langle(g-1)m\mid g\in G,m\in M\rangle$. The following lemma is useful.
@@
See [NSW, Proposition 1.6.12]. ∎
Let $E$ be a fixed elliptic curve defined over $K$. Given a finite extension $F/K$, denote by $E(F)$ (resp. $\Sh(E/F)$) the Mordell–Weil group (resp. Tate–Shafarevich group) of $E$ over $F$. Fix an algebraic closure $\overline{L}$ of $L$ and let $\operatorname{G}_{L}=\operatorname{Gal}(\overline{L}/L)$ denote the absolute Galois group of $L$. Let $\Omega_{L}$ be the set of all non-archimedean places of $L$, and for each $v\in\Omega_{L}$, let $L_{v}$ denote the completion of $L$ at $v$. Choose an algebraic closure $\overline{L}_{v}$ of $L_{v}$ and fix an embedding $\iota_{v}\mathrel{\mathop{\ordinarycolon}}\overline{L}\hookrightarrow\overline{L}_{v}$. This choice induces an embedding of Galois groups $\iota_{v}^{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{L_{v}}\hookrightarrow\operatorname{G}_{L}$, where we set $\operatorname{G}_{L_{v}}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\overline{L}_{v}/L_{v})$. For notational convenience, we write
$H^{i}(L,\cdot)\mathrel{\mathop{\ordinarycolon}}=H^{i}(\operatorname{G}_{L},\cdot),\quad H^{i}(L_{v},\cdot)\mathrel{\mathop{\ordinarycolon}}=H^{i}(\operatorname{G}_{L_{v}},\cdot)$
@@
$\kappa_{E,v}^{(n)}\mathrel{\mathop{\ordinarycolon}}\frac{E(L_{v})}{\ell^{n}E(L_{v})}\hookrightarrow H^{1}(L_{v},E[\ell^{n}])$
@@
$\kappa_{E,v}\mathrel{\mathop{\ordinarycolon}}E(L_{v})\otimes\left(\mathbb{Q}_{\ell}/\mathbb{Z}_{\ell}\right)\hookrightarrow H^{1}(L_{v},E[\ell^{\infty}]).$
@@
$\operatorname{Sel}_{\ell^{n}}(E/L)\mathrel{\mathop{\ordinarycolon}}=\ker\left(H^{1}(L,E[\ell^{n}])\to\prod_{v\in\Omega_{L}}H^{1}(L_{v},E)[\ell^{n}]\right).$
@@
$0\to E(L)\otimes\mathbb{Q}_{\ell}/\mathbb{Z}_{\ell}\to\operatorname{Sel}_{\ell^{\infty}}(E/L)\to\Sh(E/L)[\ell^{\infty}]\to 0.$
@@
$\operatorname{Sel}_{\ell}(E/L)\mathrel{\mathop{\ordinarycolon}}=\ker\left(H^{1}(L_{S}/L,E[\ell])\xrightarrow{\Phi_{S,L}}\bigoplus_{w\in S(L)}H^{1}(L_{w},E)[\ell]\right),$
@@
cf. [GCEC, p.8]. For a prime $v$ of $K$, define the restriction map
$\gamma_{v}\mathrel{\mathop{\ordinarycolon}}H^{1}(K_{v},E)[\ell]\to\bigoplus_{w\mid v}H^{1}(L_{w},E)[\ell],$
@@
$0\rightarrow E(K)\otimes\mathbb{Z}/\ell\mathbb{Z}\rightarrow\operatorname{Sel}_{\ell}(E/K)\rightarrow\Sh(E/K)[\ell]\rightarrow 0.$
@@
It follows from the assumptions above that $\gamma_{\ell}$ is injective for every prime $\ell\in S$ and thus by Proposition 3.3, $\operatorname{Sel}_{\ell}(E/L)=0$. For further details, we refer to the proof of [pathakray, Proposition 2.9]. ∎
Let $\rho_{E,\ell}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\to\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ be the Galois representation on $E[\ell]$, and let $\mathbb{Q}(E[\ell])$ be the field fixed by $\ker\rho_{E,\ell}$, so that $\operatorname{Gal}(\mathbb{Q}(E[\ell])/\mathbb{Q})\simeq\operatorname{im}\rho_{E,\ell}$. For a prime $p$ of good reduction for $E$, let $\operatorname{G}_{\mathbb{Q}_{p}}=\operatorname{Gal}(\overline{\mathbb{Q}}_{p}/\mathbb{Q}_{p})$ with inertia subgroup $\operatorname{I}_{p}$, and choose a lift $\sigma_{p}\in\operatorname{G}_{\mathbb{Q}_{p}}$ of the Frobenius. The trace and determinant of $\rho_{E,\ell}(\sigma_{p})$ satisfy
@@
For $p\in\mathfrak{T}_{E,K}$, $\widetilde{E}(k_{v})[\ell]=0$ for any of the primes of $K$ such that $v|\ell$ (cf. [pathakray, Lemma 2.11]).
@@
$\mathfrak{d}(\mathfrak{T}_{E,K})=\frac{\ell^{2}-\ell-1}{[K(\mu_{\ell})\mathrel{\mathop{\ordinarycolon}}\mathbb{Q}(\mu_{\ell})]\,(\ell^{2}-1)(\ell-1)}.$
@@ 3.2. Main results
$\varrho\mathrel{\mathop{\ordinarycolon}}\operatorname{Gal}\big(\mathbb{Q}(E[\ell])/\mathbb{Q}(\mu_{\ell})\big)\xrightarrow{\sim}\operatorname{SL}_{2}(\mathbb{F}_{\ell}).$
Set $F_{0}\mathrel{\mathop{\ordinarycolon}}=F\,\cap\,\mathbb{Q}(E[\ell])$ and
$N\mathrel{\mathop{\ordinarycolon}}=\varrho\left(\operatorname{Gal}(\mathbb{Q}(E[\ell])/F_{0})\right)\subseteq\operatorname{SL}_{2}(\mathbb{F}_{\ell}).$
@@
$\operatorname{SL}_{2}(\mathbb{F}_{\ell})\mathrel{\mathop{\ordinarycolon}}N]=[F_{0}\mathrel{\mathop{\ordinarycolon}}\mathbb{Q}(\mu_{\ell})]=\ell^{k},$
for some $1\leq\ell^{k}\leq[F\mathrel{\mathop{\ordinarycolon}}\mathbb{Q}(\mu_{\ell})]$. For primes $\ell\geq 5$, Galois [galois1846works, p. 412] showed that $\operatorname{PSL}_{2}(\mathbb{F}_{\ell})$ is simple. Therefore, $\operatorname{SL}_{2}(\mathbb{F}_{\ell})$ does not contain a proper normal subgroup $N$ with odd index $\operatorname{SL}_{2}(\mathbb{F}_{\ell})\mathrel{\mathop{\ordinarycolon}}N]$. It thus follows that $\ell^{k}=1$ and $F_{0}=\mathbb{Q}(\mu_{\ell})$. ∎
@@
For ease of notation, we set $\mathfrak{T}_{E}\mathrel{\mathop{\ordinarycolon}}=\mathfrak{T}_{E,\mathbb{Q}}$. Let $\Sigma$ be a finite set of primes containing $\ell$ and the primes at which $E$ has bad reduction. The result below is used in conjunction with Proposition 3.4 to prove Theorem 3.10, which is the main result of this section.
@@
Let $G$ be an $\ell$-group with $G=\ell^{n}$ and let $N\geq n$. There are infinitely many Galois extensions $L/\mathbb{Q}$ such that:
@@
For such a $p$, $p$ is totally ramified in $L^{p}$ and thus the inertial degree of $p$ in $L^{p}$ is $1$. Therefore, $L^{p}$ satisfies the conditions (i), (ii), (iii). We show that $p$ as above is determined by certain non-empty Chebotarev conditions. In other words, there exists a Galois extension $\mathcal{F}/\mathbb{Q}$ and a conjugation invariant non-empty subset $\mathcal{S}\subseteq\operatorname{Gal}(\mathcal{F}/\mathbb{Q})$, such that a prime $p$ which is unramified in $\mathcal{F}$ satisfies (1), (2), (3) if and only if the Frobenius element $\sigma_{p}\in\mathcal{S}$. The condition that $p\equiv 1\pmod{\ell^{N}}$ is equivalent to the requirement that $p$ splits completely in $\mathbb{Q}(\mu_{\ell^{N}})$, i.e., $\sigma_{p}=1$. The primes $q\in\Sigma$ are required to split in $L^{p}$. In other words, $q$ is an $\ell$-th power modulo $p$. Thus, $p$ is required to split completely in $\mathcal{F}^{\prime}\mathrel{\mathop{\ordinarycolon}}=\mathbb{Q}(\mu_{\ell^{N}}\cup\{\sqrt[\ell]{q}\mid q\in\Sigma\})$. On the other hand, by Lemma 3.8, $\mathcal{F}^{\prime}\cap\mathbb{Q}(E[\ell])=\mathbb{Q}(\mu_{\ell})$. Setting $\mathcal{F}\mathrel{\mathop{\ordinarycolon}}=\mathcal{F}^{\prime}\cdot\mathbb{Q}(E[\ell])=\mathbb{Q}(E[\ell]\cup\mu_{\ell^{N}}\cup\{\sqrt[\ell]{q}\mid q\in\Sigma\})$, one finds that
@@
Let $\mathcal{S}\mathrel{\mathop{\ordinarycolon}}=\mathcal{S}_{1}\times\mathcal{S}_{2}$ where $\mathcal{S}_{1}=\{1\}\subset\operatorname{Gal}(\mathcal{F}^{\prime}/\mathbb{Q}(\mu_{\ell}))$ and $\mathcal{S}_{2}\subset\operatorname{Gal}(\mathbb{Q}(E[\ell])/\mathbb{Q}(\mu_{\ell}))$ consisting of $\sigma$ such that $\operatorname{trace}\rho_{E,\ell}(\sigma)\not\equiv 2\pmod{\ell}$. It is clear that if $p\notin\Sigma$ is a prime which is unramified in $\mathcal{F}$ and $\sigma_{p}\in\mathcal{S}$, then the conditions (1)–(3) are satisfied. By the Chebotarev density theorem, this set of primes $p$ has positive density.
@@
These conditions are determined by non-empty Chebotarev sets. Condition (a) is determined in the field extension $\mathbb{Q}(E[\ell])/\mathbb{Q}(\mu_{\ell})$, and (b),(c) are determined in $\mathcal{F}^{\prime}\mathrel{\mathop{\ordinarycolon}}=L(\mu_{\ell^{N}},\sqrt[\ell]{v}\mid v\in\Sigma\cup\{p_{1},\dots,p_{m}\})$. Let $\mathcal{S}_{1}\subset\operatorname{Gal}(\mathcal{F}^{\prime}/\mathbb{Q}(\mu_{\ell}))$ be the Chebotarev set which corresponds to conditions (b) and (c), and $\mathcal{S}_{2}\subset\operatorname{Gal}(\mathbb{Q}(E[\ell])/\mathbb{Q}(\mu_{\ell}))$ corresponds to condition (a). Note that $\mathcal{F}^{\prime}$ is an $\ell$-extension of $\mathbb{Q}(\mu_{\ell})$. Lemma 3.8 thus implies that $\mathcal{F}^{\prime}\cap\mathbb{Q}(E[\ell])=\mathbb{Q}(\mu_{\ell})$. Therefore,
@@
and let $\mathcal{S}\mathrel{\mathop{\ordinarycolon}}=\mathcal{S}_{1}\times\mathcal{S}_{2}$. By construction, $\mathcal{S}\subset\operatorname{Gal}(\mathcal{F}/\mathbb{Q})$ is the non-empty Chebotarev set which determines conditions (a)–(c). Thus, by the Chebotarev density theorem, the set of primes satisfying (a)–(c) has positive density. We choose one such prime $p$ and take $p_{m+1}\mathrel{\mathop{\ordinarycolon}}=p$. Let $\chi^{(p)}\mathrel{\mathop{\ordinarycolon}}\operatorname{Gal}(\mathbb{Q}(\mu_{p})/\mathbb{Q})\rightarrow\mathbb{Z}/\ell\mathbb{Z}$ be the associated character and take $\widetilde{L}$ be the twist of $\widetilde{L}_{0}$ by $(\chi^{(p)})^{-1}$ as in Definition 2.10. This modification satisfies all the required conditions and completes the inductive argument. ∎
@@
Let $\ell\geq 5$ be a prime and $E$ be an elliptic curve over $\mathbb{Q}$. Assume that $\rho_{E,\ell}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{GL}_{2}(\mathbb{F}_{\ell})$ is surjective and that $\operatorname{Sel}_{\ell}(E/\mathbb{Q})=0$. Let $G$ be an $\ell$-group. Then there are infinitely many $G$-extensions $L/\mathbb{Q}$ for which $\operatorname{Sel}_{\ell}(E/L)=0$.